OpenStudy (anonymous):

Choose the point-slope form of the equation below that represents the line that passes through the point (−1, 6) and has a slope of −3. y − 6 = −3x − 3 y − 6 = −3(x + 1) y = −3x + 3 3x + y = 3

OpenStudy (goformit100):

i think the answer is y = −3x + 3

OpenStudy (anonymous):

a line passing through the point (a,b) and having slope m has the equation: y-b=m(x-a). so the required equation is y-6=(-3)(x+1) =>y-6=-3x-3 =>y=-3x+3